OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the point (4π, 0). y = sin(sin(x))

OpenStudy (anonymous):

y'=cosx(sin(x))*cos(x)

OpenStudy (anonymous):

4Pi = 0

OpenStudy (anonymous):

cos(sin(0))* cos(0)=1*1=1 (or so i think)

OpenStudy (anonymous):

I have to find the equation.